solve the equation by completing the square. my equation r^2-6r=0. how do i solve it?
Answer by checkley71(8403)

YES. 9 IS THE ANSWER.
--------------------------------------------------
R^2-6R=0
COMPLETING THE SQUARE REQUIRES THE (C) INTEGER BEING A PERFECT SQUARE.
WE FIND THIS SQUARE BY DIVIDING THE (B) TERM 6 BY 2 & SQUARING THE ANSWER.
6/2=3^2=9 ANSWER.
THUS THE (C) VALUE=9 ANSWER.
R^2-6R+9=0
PROOF
R^2-6R+9=(R-3)^2